OpenStudy (anonymous):

how do i write the expression 256^-3/4 in radical notation? and how do i evaluate it?

OpenStudy (anonymous):

256 = 16*16 (16)^2 *-3/4 = (16)^-3/2 16 = 4*4 (4)^2*-3/2 = 4^-3 = 1/4^3

OpenStudy (anonymous):

so this is written in radical notation: (16)^-3/2 and this is it evaluated? 1/4^3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cool. thank you!!

OpenStudy (anonymous):

that answer is WRONG!!!! it cannot be a negative because the rule of radical notation has to change it to a positive